Can you please help me solve
A line passing through (6,9)and (5,y)is parallel to a line with slope 5. What is the value of y?

Answer by venugopalramana(3286)

line passing through (6,9)and (5,y)is parallel to a line with slope 5. What is the value of y?
SLOPE OF LINE JOINING 2 POINTS (X1,Y1) AND (X2,Y2) IS (Y2-Y1)/(X2-X1)
HERE SLOPE =(Y-9)/(5-6)..SINCE PARALLEL LINES HAVE SAME SLOPE.WE HAVE
5=(Y-9)/(5-6)=-(Y-9)=-Y+9
Y=9-5=4
..